A 0.500 cm diameter plastic sphere, used in a static electricity demonstration, has a uniformly distributed 40.0 pC charge on its surface. What is the potential near its surface?

1
Expert's answer
2022-02-16T08:34:36-0500

Given:

"r=d/2=0.250\:\rm cm=2.50*10^{-3}\: m"

"q=40.0\: \rm pC=40.0*10^{-12}\:\rm C"


The potential of electric field near surface of the sphere

"V=k\frac{q}{r}=9*10^9*\frac{40.0*10^{-12}}{2.50*10^{-3}}=144\:\rm V"